# Improving IR-based Bug Localization with Semantics-Driven Query Reduction

### v1

@@ Abstract
Despite decades of research, software bug localization remains challenging due to heterogeneous content and inherent ambiguities in bug reports. Existing methods such as Information Retrieval (IR)-based approaches often attempt to match source documents to bug reports, overlooking the context and semantics of the source code. On the other hand, Large Language Models (LLM) (e.g., Transformer models) show promising results in understanding both texts and code. However, they have not been yet adapted well to localize software bugs against bug reports. They could be also data or resource-intensive. To bridge this gap, we propose, IQLoc, a novel bug localization approach that capitalizes on the strengths of both IR and LLM-based approaches. In particular, we leverage the program semantics understanding of transformer-based models to reason about the suspiciousness of code and reformulate queries during bug localization using Information Retrieval. To evaluate IQLoc, we refine the Bench4BL benchmark dataset and extend it by incorporating $\approx$30% more recent bug reports, resulting in a benchmark containing $\approx$7.5K bug reports. We evaluated IQLoc using three performance metrics and compare it against four baseline techniques. Experimental results demonstrate its superiority, achieving up to 58.52% and 60.59% in MAP, 61.49% and 64.58% in MRR, and 69.88% and 100.90% in HIT@K for the test bug reports with random and time-wise splits, respectively. Moreover, IQLoc improves MAP by 91.67% for bug reports with stack traces, 72.73% for those that include code elements, and 65.38% for those containing only descriptions in natural language. By integrating program semantic understanding into Information Retrieval, IQLoc mitigates several longstanding challenges of traditional IR-based approaches in bug localization.
@@
{highlights}
We propose a novel bug localization technique that leverages the program semantics understanding of large language models (LLMs) to identify buggy source code and reformulate search queries.
A novel technique that integrates large language models into the Information Retrieval (IR)-based bug localization and leverages their contextual reasoning.
We refined and enhanced the Bench4BL dataset by adding $\approx$30% more recent bug reports (up to September 2024), bringing the total to $\approx$7.5k bug reports.
@@ 1 Introduction
Software maintenance claimed \$2.4 trillion in 2022 from the US economy, a $\approx$84% increase over the previous two years [1]. A significant factor contributing to this increase in costs is software bug, which prevents a software system from working correctly [2]. These bugs not only hinder the functionality of software but also can lead to severe system failures (e.g., AT&T Mobility outage)[3, 4, 5, 6]. Consequently, software developers dedicate 35-50% of their time tackling these bugs [7, 8]. Zou et al. [9] recently conduct a practitioner survey where they select ten different tasks related to software bugs and collect responses from 327 software practitioners (e.g., developers, project managers, testers) affiliated with leading IT companies (e.g., Google, Meta, Microsoft, Amazon). According to their survey, 82.4% of the practitioners consider bug localization as the most important or an important task. The task involves identifying the specific locations in the software code responsible for a software bug or failure.
@@
During bug localization, developers often rely on bug reports to find the locations of problematic code. However, understanding bugs from their description in the reports and locating them in the software code is challenging and time-consuming, even for seasoned developers [10]. Bug reports contain natural language text and program artifacts such as stack traces, program elements, and code diffs [11]. The presence of these diverse contents in bug reports and the inherent ambiguities in natural language text [12] can significantly complicate the process of bug localization. To address these challenges, researchers have been actively pursuing automated solutions for bug localization over the last 50 years [13, 14, 15, 16, 17, 18, 19, 20].
@@
Over the last few decades, many methods have been employed to automatically localize software bugs. One frequently used method is Information Retrieval (IR). Existing IR-based approaches [13, 14, 15, 16] often use pre-processed texts from bug reports as search queries and attempt to match them with source code, overlooking the context or semantics of the code. This can lead to spurious matching between bug reports and source code, given the variability of natural language text describing the software bug [21, 12]. Several existing approaches incorporate code change history, version control history, or even code authoring history to improve bug localization [22, 20]. Although additional contexts of a software bug are captured, these approaches rely primarily on statistical chances rather than a comprehensive understanding of the program semantics, limiting their effectiveness in bug localization. A recent work [23] shows that these techniques perform comparably despite incorporating additional contextual information. Furthermore, another recent work [10] suggests that the performance of existing IR-based techniques might be significantly affected by their selected queries from the bug reports.
@@
Recently, several approaches have focused on making appropriate search queries based on bug reports to improve bug localization. Rahman and Roy [24] leverage both co-occurrences and syntactic dependencies among words from a bug report to capture meaningful search queries. In a recent work [11], the authors also show the use of static and hierarchical dependencies among program elements to construct queries. Other techniques [25, 26] make queries from a bug report through keyword expansion or noise removal. Although effective in making queries, these techniques could be limited by the vocabularies of bug reports, which are of varying quality [10]. This limitation restricts their capacity to comprehend contextual intricacies and locate defective source code documents. In contrast, deep learning-based techniques show promise in understanding the conceptual relationship between code and text [27, 28, 29]. Large Language Models (LLMs) advance this capability by leveraging transformer architectures and self-attention mechanisms to capture nuanced semantic relationships. Being trained on corpora beyond bug reports, these techniques can draw inferences from a larger knowledge base. However, they fail to leverage the benefits that the existing methods (e.g., IR-based) have to offer during bug localization. Moreover, they require extensive data and computational power, hindering their large-scale, sustainable adoption.
@@
To design a comprehensive solution capitalizing on the strengths of both approaches above, we propose IQLoc, a novel hybrid approach combining Information Retrieval (IR) with transformer-based models to localize software bugs. Textual relevance alone may broaden a search space, potentially delivering unrelated results. Our approach narrows down the search space by leveraging program semantics understanding of the pre-trained language models during bug localization. First, IQLoc retrieves the top K source documents textually relevant to a bug report (a.k.a., query) using an IR-based approach (e.g., the BM25 algorithm [30]). Second, it trains a transformer-based, cross-encoder model to assess their relevance (based on their program semantics) to the bug report and narrows down the search space. Finally, IQLoc reduces the search query leveraging the above documents, reranks the documents using the query, and returns the suspicious source code documents.
@@
We selected an existing benchmark dataset – Bench4BL [23] – for our experiments and extended it with recent bug reports (i.e., submitted until September 2024). To refine the dataset, we selected the bug reports that include version information and have corresponding relevant documents in their respective GitHub repositories. We further expanded it with $\approx$30% more recent bug reports, resulting in a final set of 7,483 bug reports. To evaluate our proposed technique, we employed three appropriate and widely used metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. We compare our technique with four appropriate baselines from literature – BLUiR [19], Blizzard [11], DNNLoc [17] and RLocator [31]. Across various measures, IQLoc consistently outperformed these techniques, with improvements of up to 58.52% and 60.59% in MAP, 61.49% and 64.58% in MRR, and 69.88% and 100.90% in HIT@K for the test bug reports with random and time-wise splits, respectively. Additionally, IQLoc improves bug localization over baselines in terms of MAP and MRR, achieving 91.67% and 86.49% for bug reports with stack traces, 72.73% and 66.67% for those containing program artifacts, and 65.38% and 64.29% for natural language-only bug reports.
@@
  • A comprehensive evaluation of IQLoc using three popular metrics, two different splits of the dataset, and comparison with four baseline techniques [19, 11, 17, 31] from the literature.
@@
Table 1: An Example of Bug Report and Search Queries
Bug ID # 1416 (Spring Webflow) Rank
Bug Title Turn off snapshot creation when max-snapshots=0 72
Bug Description Creating snapshots involves serializing and compressing the flow execution object, which can lead to issues with some PersistenceContext providers. Having the ability to turn off snapshot creation will provide an option when such issues occur at the cost of losing back button support. 57
Baseline query Bug Title + Bug Description 53
IQLoc persistencecontext snapshot involves creation losing support flow object turn execution ability providers serializing 1
@@ 2 Motivational Example
In this section, we demonstrate the effectiveness of our approach –IQLoc– in bug localization using a motivating example. We present an example where our approach outperforms the baseline methods. Table 1 shows an example bug report from Spring Workflow (SFW) and Fig. 1 shows the associated buggy code. When the pre-processed version of the bug report is executed as a query by an IR-based search engine (e.g., Elasticsearch), the first buggy document is found at the 53rd position, which is not an ideal result. As shown in Table 1, the pre-processed version of the title or description fields also does not make a good query.
The example bug report discusses an issue involving snapshot creation, which could affect several classes including PersistentContext. Interestingly, the root cause of the bug involves the serialization/compression of the FlowExecution object. However, traditional techniques such as BLUiR and Blizzard consistently select PersistenceContext as a query keyword since it is found in the bug report. This leads to poor retrieval performance and they retrieve their first buggy documents at the 13th and 18th positions, respectively. In contrast, the seminal work on deep learning-based bug localization –DNNLoc– considers multiple features (e.g., textual relevance, change history, collaborative filtering), but performs poorly by retrieving the buggy document at the 28th position.
On the other hand, IQLoc attempts to understand the context of a reported bug from the source code. In particular, its cross-encoder module learns to connect bug reports to their corresponding buggy code through transformer-based, contextual learning. The cross-encoder module of our technique was able to rank the buggy code at the 11th position. Our module achieves that by capturing the context of the bug from the bug report and making a connection with the intent of the buggy code (e.g., snapshot creation from FlowExecution). More interestingly, upon reformulating the query (a.k.a., preprocessed bug report) leveraging the IR-based and cross-encoder-based components, our technique delivers the buggy source document (e.g., Fig 1) at the 1st position, which is ideal and highly promising.
@@ 3.1 Program Semantics and Context
Program semantics refer to the meaning or behavior of a computer program, explaining how it processes inputs, performs computations, and produces outputs [32]. They provide a structured way to reason about a program’s functionality, correctness, and execution behavior [33]. However, program semantics cannot be fully understood in isolation — they require context. Variables, functions, and control flows must be interpreted by capturing their surrounding program elements. Existing literature [34] suggests that context is not a static property but rather something that emerges dynamically from the interaction and relationships among program elements. To localize software bugs, we need to analyze not only the source code but also its semantics within its context. In particular, we leverage the Large Language Models’ ability to understand program semantics in localizing software bugs.
@@ 3.2 Cross-Encoders
The cross-encoder [35] is a transformer-based approach used for reranking documents by enabling deep interactions between the input context and the candidate. This approach processes the context and the candidate jointly within a single transformer model. The inputs are concatenated with a special separator token (e.g., [SEP]) and passed through the transformer, allowing self-attention to capture fine-grained dependencies between them (Fig. 2).
@@
Prior studies [36, 37] have demonstrated the cross-encoder’s advantages in refining retrieved results by analyzing relationships between queries and documents. We used a cross-encoder in our first study, IQLoc, to rerank source documents based on program semantic relevance between source code and bug reports.
@@ 4.1 Fine-tune Cross-Encoder Model
In IQLoc, we employ a cross-encoder model to determine relevance between a bug report and a code segment (Step 3, Fig. 4). First, we fine-tune the base model of CodeBERT, a transformer-based encoder model [38], using bug-fix changes and enable it to differentiate between buggy and non-buggy instances during inference (Fig. 3a). In particular, we append feed-forward network to the pre-trained CodeBERT model to achieve classification. Due to its pre-training on a large amount of natural language text and source code, the model is well-suited for our task involving bug reports and source code documents.
To fine-tune the model, we first parse each source code document from the training corpus and extract its buggy methods by leveraging the bug-fix diffs from corresponding version control history (Step 1c, Fig. 3a). We detailed this method extraction process in Section 5.1. We also establish connection between bug reports and corresponding buggy code using appropriate heuristics [23]. Then we treat bug reports (Step 1a) and buggy code pairs (Step 1e) as positive samples in our training dataset. These pairs are fed into the cross-encoder model for training with the goal of establishing positive contextual associations between buggy methods and their respective bug reports (Step 2, Fig. 3a).
To train the model to differentiate between buggy and non-buggy samples, we also add negative pairs where each pair consists of a bug report (Step 1a, Fig. 3a) and a randomly selected method body (Step 1b-dashed) to distinguish buggy from non-buggy code by leveraging program semantics. The capability stems from CodeBERT’s pre-training on a large bimodal corpus, which equips it with rich semantic representations of code structure and intent [38], as evidenced by its performance on downstream tasks (e.g., code summarization, clone detection [39, 40, 41]). Fine-tuning further adapts these representations by reinforcing links between real-world bug reports and their corresponding buggy code, enabling the model to capture subtle semantic cues associated with buggy behavior.

### v2

@@ Abstract
Despite decades of research, software bug localization remains challenging due to heterogeneous content and inherent ambiguities in bug reports. Existing methods, such as Information Retrieval (IR)-based approaches, often attempt to match source documents to bug reports, overlooking the context and semantics of the source code. On the other hand, Large Language Models (LLMs) (e.g., Transformer models) show promising results in understanding both texts and code. However, they have not yet been adapted well to localize software bugs using bug reports. They could also be data or resource-intensive. To bridge this gap, we propose, IQLoc, a novel approach that capitalizes on the strengths of both IR and LLMs for bug localization. In particular, we leverage the transformer-based model’s understanding of code semantics to reason about its suspiciousness and to reformulate search queries and thus enhance bug localization using Information Retrieval. To evaluate IQLoc, we refine the Bench4BL benchmark dataset and extend it by incorporating $\approx$30% more recent bug reports, resulting in a benchmark containing $\approx$7.5K bug reports. We evaluated IQLoc using three performance metrics and compare it against eight baseline techniques. Experimental results demonstrate its superiority, achieving up to 100.40% and 78.08% in MAP, 61.49% and 64.58% in MRR, and 76.98% and 100.90% in HIT@K for the test bug reports with random and time-wise splits, respectively. Moreover, IQLoc improves MAP by 118.70% for bug reports with stack traces, 111.87% for those that include code elements, and 127.45% for those containing only descriptions in natural language. By integrating program semantic understanding into Information Retrieval, IQLoc mitigates several longstanding challenges of traditional IR-based approaches in bug localization.
@@ 1 Introduction
Software maintenance claimed \$2.4 trillion in 2022 from the US economy, a $\approx$84% increase over the previous two years [US_maintenance_stats_2022]. A significant factor contributing to this increase in costs is software bug, which prevents a software system from working correctly [SE_Terminology]. These bugs not only hinder the functionality of software but also can lead to severe system failures (e.g., AT&T Mobility outage)[FCC2024_ATnT, Reuters2024_ATnT, Summary_ATnT, nasaMariner]. Consequently, software developers dedicate 35-50% of their time tackling these bugs [o2017debugging_stats, britton2013reversible_stats]. Zou et al. [practitioners_bug_localization_study] recently conduct a practitioner survey where they select ten different tasks related to software bugs and collect responses from 327 software practitioners (e.g., developers, project managers, testers) affiliated with leading IT companies (e.g., Google, Meta, Microsoft, Amazon). According to their survey, 82.4% of the practitioners consider bug localization as the most important or an important task. The task involves identifying the specific locations in the software code responsible for a software bug or failure.
@@
During bug localization, developers often rely on bug reports to find the locations of problematic code. However, understanding bugs from their description in the reports and locating them in the software code is challenging and time-consuming, even for seasoned developers [rahman2021forgotten]. Bug reports contain natural language text and program artifacts such as stack traces, program elements, and code diffs [blizzard]. The presence of these diverse contents in bug reports and the inherent ambiguities in natural language text [furnas1987vocabulary] can significantly complicate the process of bug localization. To address these challenges, researchers have been actively pursuing automated solutions for bug localization over the last 50 years [ir_bug_localization1, ir_bug_localization2_spectra, ir_bug_localization3_topic, ir_bug_localization4, dnnloc_ir, ir_localization_lda_buglocator, saha_bleuir, ir_ml_amalgam].
@@
Over the last few decades, many methods have been employed to automatically localize software bugs. One frequently used method is Information Retrieval (IR). Existing IR-based approaches [ir_bug_localization1, ir_bug_localization2_spectra, ir_bug_localization3_topic, ir_bug_localization4] often use pre-processed texts from bug reports as search queries and attempt to match them with source code, overlooking the context or semantics of the code. This can lead to spurious matching between bug reports and source code, given the variability of natural language text describing the software bug [query_keyword, furnas1987vocabulary]. Several existing approaches incorporate code change history, version control history, or even code authoring history to improve bug localization [bl_code_change_BLIA, ir_ml_amalgam]. Although additional contexts of a software bug are captured, these approaches rely primarily on statistical chances rather than a comprehensive understanding of the program semantics, limiting their effectiveness in bug localization. A recent work [bench4bl] shows that these techniques perform comparably despite incorporating additional contextual information. Furthermore, another recent work [rahman2021forgotten] suggests that the performance of existing IR-based techniques might be significantly affected by their selected queries from the bug reports.
@@
Recently, several approaches have focused on making appropriate search queries based on bug reports to improve bug localization. Rahman and Roy [strict] leverage both co-occurrences and syntactic dependencies among words from a bug report to capture meaningful search queries. In a recent work [blizzard], the authors also show the use of static and hierarchical dependencies among program elements to construct queries. Other techniques [chaparro2017_ob_eb, sisman2012incorporating] make queries from a bug report through keyword expansion or noise removal. Although effective in making queries, these techniques could be limited by the vocabularies of bug reports, which are of varying quality [rahman2021forgotten]. This limitation restricts their capacity to comprehend contextual intricacies and locate defective source code documents. In contrast, deep learning-based techniques show promise in understanding the conceptual relationship between code and text [cao2020bugpecker, zhu2020cooba, xiao2023bugradar]. Large Language Models (LLMs) advance this capability by leveraging transformer architectures and self-attention mechanisms to capture nuanced semantic relationships. Being trained on corpora beyond bug reports, these techniques can draw inferences from a larger knowledge base. However, they fail to leverage the benefits that the existing methods (e.g., IR-based) have to offer during bug localization. Moreover, they require extensive data and computational power, hindering their large-scale, sustainable adoption.
@@
To design a comprehensive solution capitalizing on the strengths of both approaches above, we propose IQLoc, a novel hybrid approach combining Information Retrieval (IR) with transformer-based models to localize software bugs. Textual relevance alone may broaden a search space, potentially delivering unrelated results. Our approach narrows down the search space by leveraging program semantics understanding of the pre-trained language models during bug localization. First, IQLoc retrieves the top K source documents textually relevant to a bug report (a.k.a., query) using an IR-based approach (e.g., the BM25 algorithm [robertson1995okapi_bm25]). Second, it trains a transformer-based, cross-encoder model to assess their relevance (based on their program semantics) to the bug report and narrows down the search space. Finally, IQLoc reduces the search query leveraging the above documents, reranks the documents using the query, and returns the suspicious source code documents.
@@
We selected an existing benchmark dataset – Bench4BL [bench4bl] – for our experiments and extended it with recent bug reports (i.e., submitted until September 2024). To refine the dataset, we selected the bug reports that include version information and have corresponding relevant documents in their respective GitHub repositories. We further expanded it with $\approx$30% more recent bug reports, resulting in a final set of 7,483 bug reports. To evaluate our proposed technique, we employed three appropriate and widely used metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. We compare our technique with eight appropriate baselines including – BLUiR [saha_bleuir], Blizzard [blizzard], BRTracer [BRTracer], BLIA [bl_code_change_BLIA], LRBL [LRBL_learning2rank], LLmiRQ [LLmiRQ], DNNLoc [dnnloc_ir] and RLocator [RLocator]. Across various measures, IQLoc consistently outperformed these techniques, with improvements of up to 100.40% and 78.08% in MAP, 61.49% and 64.58% in MRR, and 76.98% and 100.90% in HIT@K for the test bug reports with random and time-wise splits, respectively. Additionally, IQLoc improves bug localization over baselines in terms of MAP and MRR, achieving 118.70% and 86.49% for bug reports with stack traces, 111.87% and 71.09% for those containing program artifacts, and 127.45% and 75.23% for natural language-only bug reports.
@@
  • A comprehensive evaluation of IQLoc using three popular metrics, two different splits of the dataset, and comparison with eight baseline techniques [saha_bleuir, blizzard, BRTracer, bl_code_change_BLIA, LRBL_learning2rank, LLmiRQ, dnnloc_ir, RLocator] from the literature.
@@
Table 1: A Motivational Example Illustrating Bug Localization Techniques and Approaches (Spring Webflow # SWF-1416)
Title: Turn off snapshot creation when max-snapshots = 0
Description: Creating snapshots involves serializing and compressing the flow execution object, which can lead to issues with some PersistenceContext providers. Having the ability to turn off snapshot creation will provide an option when such issues occur at the cost of losing back button support.
Technique Approach Rank
Bug Title (Query) Traditional IR (Bug Title Only) 72
Bug Description (Query) Traditional IR (Bug Description Only) 57
Baseline Query Traditional IR (Bug Title + Bug Description) 53
Blizzard [blizzard] Traditional IR (Context-Aware Query Reformulation) 18
LRBL [LRBL_learning2rank] Learning-to-Rank (Feature-Based Ranking) 19
DNNLoc [dnnloc_ir] Deep Learning (Neural Ranking Model) 28
RLocator [RLocator] Reinforcement Learning (Ranking Optimization) 6
IQLoc (Proposed) Program Semantic Understanding 1
@@ 2 Motivational Example
In this section, we demonstrate the effectiveness of our approach –IQLoc– in bug localization using a motivating example. We present an example where our approach outperforms the baseline methods. Table 1 shows an example bug report (Bug ID#SWF-1416) from Spring Webflow (SFW) [springwebflow] and Fig. 1 shows the associated buggy code. When the pre-processed version of the bug report is executed as a query by an IR-based search engine (e.g., Elasticsearch), the first buggy document is found at the 53rd position, which is not an ideal result. As shown in Table 1, the pre-processed version of the title or description field also does not make a good query.
The example bug report discusses an issue involving snapshot creation, which could affect several classes including PersistentContext. Interestingly, the root cause of the bug involves the serialization/compression of the FlowExecution object. However, traditional techniques such as BLUiR and Blizzard consistently select PersistenceContext as a query keyword since it is found in the bug report. This leads to poor retrieval performance and they retrieve their first buggy documents at the 13th and 18th positions, respectively.
In contrast, LRBL, a learning-to-rank technique, leverages a mix of heterogeneous features including lexical similarity, class-name overlap, collaborative filtering, and historical bug-fixing information to rank suspicious source documents. Similarly, DNNLoc, the seminal work on deep learning–based bug localization, also considers multiple features (e.g., textual relevance, change history, and collaborative filtering) and employs a neural ranking model to predict suspiciousness scores. Nevertheless, LRBL and DNNLoc perform poorly for this bug, retrieving the buggy document at the 19th and 28th positions, respectively. On the other hand, RLocator, a reinforcement learning–based technique, formulates bug localization as a ranking optimization problem. However, it still fails to capture the nuances of program understanding and ranks the buggy document at the 6th position, which is not ideal.
In comparison, IQLoc attempts to understand the context of a reported bug from the source code. In particular, its cross-encoder module learns to connect bug reports to their corresponding buggy code through transformer-based, contextual learning. The cross-encoder module of our technique was able to rank the buggy code at the 11th position. Our module achieves that by capturing the context of the bug from the bug report and making a connection with the intent of the buggy code (e.g., snapshot creation from FlowExecution). More interestingly, upon reformulating the query (a.k.a., preprocessed bug report) leveraging the IR-based and cross-encoder-based components, our technique delivers the buggy source document (e.g., Fig. 1) at the 1st position, which is ideal and highly promising.
@@ 3.1 Program Semantics and Context
Program semantics refer to the meaning or behavior of a computer program, explaining how it processes inputs, performs computations, and produces outputs [semantics_programming_lang]. They provide a structured way to reason about a program’s functionality, correctness, and execution behavior [semcoder_program_semantics]. However, program semantics cannot be fully understood in isolation — they require context. Variables, functions, and control flows must be interpreted by capturing their surrounding program elements. Existing literature [context_definition] suggests that context is not a static property but rather something that emerges dynamically from the interaction and relationships among program elements. To localize software bugs, we need to analyze not only the source code but also its semantics within its context. In particular, we leverage the Large Language Models’ ability to understand program semantics in localizing software bugs.
@@ 3.2 Cross-Encoders
The cross-encoder [polyencoder] is a transformer-based approach used for reranking documents by enabling deep interactions between the input context and the candidate. This approach processes the context and the candidate jointly within a single transformer model. The inputs are concatenated with a special separator token (e.g., [SEP]) and passed through the transformer, allowing the self-attention mechanism to capture fine-grained dependencies between them (Fig. 2).
@@
Prior studies [crossencoder_example_1, crossencoder_example_2] have demonstrated the cross-encoder’s advantages in refining retrieved results by analyzing relationships between queries and documents. Therefore, we use a cross-encoder in IQLoc to rerank source documents based on program semantic relevance between source code and bug reports.
4 Methodology
@@ 4.1 Fine-tune Cross-Encoder Model
In IQLoc, we employ a cross-encoder model to determine relevance between a bug report and a code segment (Step 3, Fig. 4). First, we fine-tune the base model of CodeBERT, a transformer-based encoder model [feng2020codebert], using bug-fix changes and enable it to differentiate between buggy and non-buggy instances during inference (Fig. 3(a)). In particular, we append feed-forward network to the pre-trained CodeBERT model to achieve classification. Due to its pre-training on a large amount of natural language text and source code, the model is well-suited for our task involving bug reports and source code documents.
To fine-tune the model, we first parse each source code document from the training corpus and extract its buggy methods by leveraging the bug-fix diffs from corresponding version control history (Step 1c, Fig. 3(a)). We detailed this method extraction process in Section 5.1. We also establish connection between bug reports and corresponding buggy code using appropriate heuristics [bench4bl]. Then we treat bug reports (Step 1a) and buggy code pairs (Step 1e) as positive samples in our training dataset. These pairs are fed into the cross-encoder model for training with the goal of establishing positive contextual associations between buggy methods and their respective bug reports (Step 2, Fig. 3(a)).
To train the model to differentiate between buggy and non-buggy samples, we also add negative pairs where each pair consists of a bug report (Step 1a, Fig. 3(a)) and a randomly selected method body (Step 1b-dashed) to distinguish buggy from non-buggy code by leveraging program semantics. The capability stems from CodeBERT’s pre-training on a large bimodal corpus, which equips it with rich semantic representations of code structure and intent [feng2020codebert], as evidenced by its performance on downstream tasks (e.g., code summarization, clone detection [codeClone1CodeBert, codeClone2ProbeCodeBert, programTranslationCodeBert]). Fine-tuning further adapts these representations by reinforcing links between real-world bug reports and their corresponding buggy code, enabling the model to capture subtle semantic cues associated with buggy behavior.
To fine-tune our cross-encoder model, we configure the hyperparameters following standard practices [feng2020codebert]. We set our initial learning rate to $1\times 10^{-4}$ and optimized the model automatically using PyTorch’s ReduceLROnPlateau scheduler [pytorch], which halved the rate whenever the validation loss did not improve for consecutive epochs (e.g., 2). The batch size was set to 8 to balance memory usage and gradient stability, given the computational constraints of our cloud server. We set the number of epochs at 100 at the beginning of the tuning; however, the effective training duration was determined dynamically, as the scheduler progressively reduced the learning rate and training stabilized once the loss plateaued. This configuration aligns with widely adopted fine-tuning strategies in the literature, which emphasize adaptive scheduling and validation-based stopping to ensure stable convergence and prevent overfitting [devlin2019bert_HypParam].